Assembled most of the connectors on the new (green) manifolds. Two blanking pulgs stuck when mounting - cannot remove them or tighten them. Swarf? Will ask Davide's help tomorrow. Cannot use the manifold until this is fixed.
Checked FEE assembly in cassette. Need to shorten black Delrin supports by ~6 mm. Holes should end up in the correct position thanks to rails being effectively longer (same length, different hole position on FEE cassette). Will ask Davide to skim one if possible. If not, will have to mount without supports.
Finally managed to get the CAEN program working to control N1419 remotely from Baking Pi. Plotting bias and leakage current overnight to test stability. See attached. IE514 sensor off for these tests. |